# InternVideo2: Scaling Video Foundation Models for Multimodal Video Understanding

### v1

@@ Abstract
We introduce InternVideo2, a new video foundation model (ViFM) that achieves the state-of-the-art performance in action recognition, video-text tasks, and video-centric dialogue. Our approach employs a progressive training paradigm that unifies the different self- or weakly-supervised learning frameworks of masked video token reconstruction, cross-modal contrastive learning, and next token prediction. Different training stages would guide our model to capture different levels of structure and semantic information through different pretext tasks. At the data level, we prioritize the spatiotemporal consistency by semantically segmenting videos and generating video-audio-speech captions. This improves the alignment between video and text. We scale both data and model size for our InternVideo2. Through extensive experiments, we validate our designs and demonstrate the state-of-the-art performance on over 60 video and audio tasks. Notably, our model outperforms others on various video-related captioning, dialogue, and long video understanding benchmarks, highlighting its ability to reason and comprehend long temporal contexts.
@@ 1 Introduction
The pursuit of transferable spatiotemporal representations is a critical research area within vision understanding, holding diverse applications across domains such as video searching Gabeur et al. (2020), game control Bruce et al. (2024), robotic learning Driess et al. (2023), self-driving Zablocki et al. (2022), and scientific studies Team et al. (2023). Recently, the advancement of Large Language Models (LLMs) Brown et al. (2020); OpenAI (2023a); Touvron et al. (2023a, b) and their multimodal variations (MLLMs) OpenAI (2023b); Gong et al. (2023); Liu et al. (2023); Zhu et al. (2023a) has had a profound impact on vision research and other disciplines, thanks to their learned world models OpenAI (2023c); Yang et al. (2024). Embedding videos effectively into these world models and harnessing their capabilities to enhance video embedding have emerged as pivotal tasks for advancing video understanding.
@@
Previous research has identified several effective learning schemes for video representation, including reconstructing videos with masked inputs (MAE He et al. (2022) and VideoMAE Wang et al. (2023a)), aligning videos with languages Xu et al. (2021); Yan et al. (2022); Li et al. (2023a), and predicting the next token using videos Alayrac et al. (2022); Sun et al. (2023a); Li et al. (2023b). These approaches have been shown to be complementary and can be unified through a progressive training scheme. Notably, methods such as UMT Li et al. (2023a), InternVideo Wang et al. (2022), and VideoPrism Zhao et al. (2024) have utilized a two-stage training approach involving masked reconstruction and multimodal contrastive learning, leading to improved performance in downstream tasks. These studies have demonstrated the benefits of joint learning for enhancing spatiotemporal perception and video-language alignment. This paper extends upon this scheme by incorporating video-based next token prediction and scaling the entire training process, including models and data.
@@
The proposed video foundation model InternVideo2, is built through a progressive learning scheme. The learning involves three stages aimed at improving the model’s spatiotemporal perception, aligning it with semantics from other modalities, and enhancing its world modeling through next token prediction Alayrac et al. (2022). In the initial stage of masked video token reconstruction, the model learns to reconstruct masked video tokens, allowing the video encoder to develop basic spatiotemporal perception. To estimate the masked tokens, vision encoders (InternViT Chen et al. (2023a) and VideoMAE-g Wang et al. (2023a)) trained differently are employed as proxies. In the next stage of multimodal learning, the architecture is expanded to include audio and text encoders. This not only improves the alignment between videos and text but also grants InternVideo2 the ability to handle video-audio tasks. By incorporating these additional modalities, the model’s understanding of videos is enriched and aligned with the semantics provided by audios. Finally, in the next-token prediction stage, a video-centric dialogue system and the corresponding instruction-finetuning dataset are leveraged to train InternVideo2. This transfer learning process allows the model to benefit from the knowledge acquired by LLM and others. By connecting InternVideo2 to LLMs, the video encoder is further updated through next-token prediction training, enhancing its ability to generate contextually appropriate next tokens.
@@
In pretraining, we emphasize incorporating spatiotemporal consistency in the data. We utilize a large-scale multimodal video-centric dataset consisting of 412M data entries, which includes 2M videos, 50M video-text pairs, 60M video-audio-speech-text pairs, and 300M image-text pairs. In multimodal pretraining data, we segment videos into clips semantically, and we focus on recalibrating the descriptions with videos using three modalities: audio, video, and speech. To this end, we generate captions for these three modalities separately. Then individual captions are fused together to create a more comprehensive description. By incorporating audio, video, and speech modalities in the training process, we provide a richer and more detailed understanding of the videos. This holistic data approach to video-text alignment improves the model’s ability to comprehend and interpret the video accurately.
@@
We evaluate InternVideo2 across a range of video-related tasks. These tasks span from basic spatiotemporal perception, such as action recognition, to more complex reasoning tasks, such as long video or procedure-aware question-answering (QA), as given in Fig. 1. The results (in Sec. 5) demonstrate that InternVideo2 achieves state-of-the-art performance on multiple tasks, able to analyzes and reasons sequences of actions. These highlight that InternVideo2 excels not only in video perception and video-language alignment but also in modeling the world. Its top performance on various benchmarks in the field of multimodal language models (MLLM) signifies its capability to effectively capture and understand video content. These empirical findings validate that InternVideo2 serves as a qualified video encoder for future exploration in video understanding.
In summary, our contributions to video understanding are as follows.
@@
This paper introduces InternVideo2, a competitive video foundation model that leverages masked reconstruction, crossmodal contrastive learning, and next token prediction to make model perceptive, semantic, and capable of reasoning in video understanding. It achieves the state-of-the-art performance in 65 out of 74 video / audio tasks. Our model demonstrates superior performance in video-related dialogue and long video understanding, highlighting its potential in various world model studies and applications.
@@
We provide an enhanced dataset to train InternVideo2. This includes the validation and incorporation of audio data during training, as well as the implementation of annotation techniques. These improvements result in significant enhancements in model generalization and performance.
@@ 2 Related Work
Video Understanding. Studies on learning video foundation models becomes increasingly crucial considering its wide applications Li and Wang (2020); Xu et al. (2021); Li et al. (2023a, 2022a); Hu et al. (2022); Dou et al. (2022); Shen et al. (2021); Yao et al. (2021); Sun et al. (2019); Zhu and Yang (2020); Wang et al. (2022); Chen et al. (2022a); Zellers et al. (2021, 2022); Zeng et al. (2023a, b); Chen et al. (2023b, c); He et al. (2023); Chen et al. (2024a); Zhao et al. (2024); Wang et al. (2023b); Yan et al. (2022); Feichtenhofer et al. (2022); Wang et al. (2023c); Miech et al. (2020); Tong et al. (2022); Wang et al. (2023a); Li et al. (2023c). Typical methods in building video foundation models (ViFM) include video-text contrastive learning Xu et al. (2021); Wang et al. (2023c, 2022), masked video modeling Tong et al. (2022); Wang et al. (2023a, 2022); Fu et al. (2021), and next token prediction Alayrac et al. (2022); Sun et al. (2023a, b). Notably, All-in-one Wang et al. (2023c) utilizes a single backbone with unified multiple pretraining objectives. On the other hand, UMT Li et al. (2023a) combines masked modeling with video-text contrastive learning, demonstrating strong performance in both action recognition and video-language tasks. Another approach is mPLUG-2 Xu et al. (2023), which introduces a new design for modulating different modalities. It shares a common module across modalities to enhance relations while incorporating modality-specific modules for discrimination. In addition to video-text pretraining, researchers have also explored the use of audio information in videos to improve performance. MERLOT Reserve Zellers et al. (2022) learns video representations using a large-scale dataset of video-speech-transcript pairs. VALOR Chen et al. (2023c) employs independent encoders for video, audio, and text and trains a joint visual-audio-text representation. VAST Chen et al. (2024a) constructs an audio-visual-speech dataset and develops a multimodal backbone that excels in video-audio-related tasks. VideoPrism Zhao et al. (2024) combines video-text contrastive learning and video token reconstruction on a combination of public and proprietary videos, achieving leading results across various video tasks.
@@
Multimodal Large Language Models. With advances in large language models (LLMs) Devlin et al. (2018); Raffel et al. (2020); Brown et al. (2020); Wei et al. (2021); Chowdhery et al. (2022); Sun et al. (2023a, b), their multimodal versions (MLLMs) gain a lot heat considering its open-world visual task execution by simple instructions. Seminal works like Flamingo Alayrac et al. (2022) shows outstanding zero/few-shots performances over a range of multimodal tasks Goyal et al. (2017a); Plummer et al. (2015); Xu et al. (2016); Marino et al. (2019). Public MLLMs Zhu et al. (2023b); Liu et al. (2023); Gong et al. (2023) e.g. LLaVA Liu et al. (2023) and InstructBLIP Dai et al. (2023) give visual instruction-tuning data for dialogue experience. Some video-related MLLM (VideoLLM) VideoChatGPT Maaz et al. (2023a) and Valley Luo et al. (2023) utilize ChatGPT to generate video instruction-tuning data, aiming to enhance instruction-following capabilities for real-world video comprehension.
@@ 3 Method
We learn InternVideo2 in three stages, illustrated in Fig. 2. The stages include spatiotemporal token reconstruction, video-audio-speech-language contrastive learning, and joint training of a large language model (LLM).
@@
The video encoder used in InternVideo2 follows the Vision Transformer (ViT) Dosovitskiy et al. (2020) and includes additional projection layers for distillation. Inspired by previous works Chen et al. (2023a); Yu et al. (2022), we introduce attention pooling to the ViT. For input videos, we sparsely sample 8 frames and perform a 1414 () spatial downsampling. These spatiotemporal tokens are then concatenated with a class token and combined with 3D position embeddings. The hyperparameters used for ViT-1B and -6B are the same as CoCa-1B Yu et al. (2022) and InternViT-6B Chen et al. (2023a), respectively. The details of ViT-6B architecture are given in Supp. Tab. 21.
3.1 Stage1: Reconstructing Masked Video Tokens
We exploit two expert models to guide the video encoder to conduct masked token-level reconstruction. Specifically, we adopt InternVL-6B Chen et al. (2023a) and VideoMAEv2-g Wang et al. (2023a) to transfer unmasked knowledge via simple projection layers. When training, we input the full videos into different teachers and masked out 80% of the tokens frame by frame, under the semantic guidance of the multimodal model InternVL and motion-aware model VideoMAEv2. We only align the related unmasked tokens, by minimizing their mean squared error (MSE). The learning objective of this stage is formed by reconstructing the remaining tokens as:
where , , and are our video encoder, InternViT-6B Chen et al. (2023a), and ViT-g of VideoMAEv2, respectively. stands for the token index and is the corresponding token extracted from InternVideo2 of input video . is the normalization factor. and balance the influence between the employed models.
In implementation, we align the last layers of InternVL, last layers of VideoMAE, and the final output token of InternVL. Different losses are simply summed for optimization. After pretraining, we drop those projection layers and only use the basic encoder. Compared with only using the multimodal model in UMT and VideoPrism, our strategy makes the vision encoder multimodal-friendly as well as enhances its temporal sensitivity for action modeling.
@@ 3.2 Stage 2: Aligning Video to Audio-Speech-Text
We exploit the correspondence between video and audio, speech, and text to align InternVideo2 to semantics explicitly. In structure, though InternVideo2 has a huge video encoder, its employed audio and text encoders are relatively lightweight. The used audio encoder is a 12-layer transformer initialized with BEATs Chen et al. (2023d) (90M). It takes in audio features, which are 64-dimensional log Mel filterbank spectrograms using a 25ms Hamming window, transformed from 10-second-long clips, padding with zeros. For the text and speech encoders, we initialize the text encoder and multimodal decoder using Bert-Large Devlin et al. (2018). Specifically, we utilize the initial 19 layers of Bert-Large as the text encoder, with the subsequent 5 layers equipped with cross-attention layers serving as the multimodal decoder.
For pretraining objectives, we establish alignment and fusion across modalities via text, including video, audio, image, and speech. We employ crossmodal contrastive and matching losses with masked language reconstruction one as:
where , , and are hyperparameters to tradeoff different losses. The employed and are standard and referenced from Cheng et al. (2022). Specifically, the crossmodal contrastive learning is given as:
@@
where computes the matching likelihood between video and text , respectively. denotes whether the given video and text are paired () or not ().
@@
For training efficiency, we employ masking learning, aligning unmasked video tokens to tokens from other modalities first, then using full video tokens reconstruction shortly. Specifically, it consists of two steps as follows:
@@
We freeze the audio encoder, focusing primarily on aligning visual and text features. We employ the complete set of image, video, and audio-video data for pre-training, where in this training (the combinations denote the concatenated features from the used modalities).
@@
We freeze the vision encoder to achieve joint alignment of audio, visual, and text features. We conduct post-pretraining using a small subset of image and video data (25M), along with the complete set of audio (0.5M) and audio-video data (50M). Given that the parameters of the largest ViT-6B model are frozen, we abandon masking strategies here for the consistency with the inference, minimizing the degradations on downstream performance brought by masking. Here .

### v2

# InternVideo2: Scaling Foundation Models for Multimodal Video Understanding

@@ Abstract
We introduce InternVideo2, a new family of video foundation models (ViFM) that achieve the state-of-the-art results in video recognition, video-text tasks, and video-centric dialogue. Our core design is a progressive training approach that unifies the masked video modeling, crossmodal contrastive learning, and next token prediction, scaling up the video encoder size to 6B parameters. At the data level, we prioritize spatiotemporal consistency by semantically segmenting videos and generating video-audio-speech captions. This improves the alignment between video and text. Through extensive experiments, we validate our designs and demonstrate superior performance on over 60 video and audio tasks. Notably, our model outperforms others on various video-related dialogue and long video understanding benchmarks, highlighting its ability to reason and comprehend longer contexts.
@@ 1 Introduction
Learning transferrable spatiotemporal representations is a critical research area in computer vision, holding diverse applications across domains such as video search Gabeur et al. (2020), game control Bruce et al. (2024), robotic learning Driess et al. (2023), self-driving Zablocki et al. (2022), and scientific studies Team et al. (2023). Recently, the advancement of Large Language Models (LLMs) Brown et al. (2020); OpenAI (2023a); Touvron et al. (2023a, b) and their multimodal variations (MLLMs) OpenAI (2023b); Gong et al. (2023); Liu et al. (2023); Team et al. (2023) have had a profound impact on vision research and other disciplines. Embedding videos effectively into these large models and harnessing their capabilities to enhance video understanding performance has emerged as pivotal tasks Li et al. (2023c); Maaz et al. (2023b).
@@
Previous research has identified several effective learning schemes for video representations, including reconstructing videos with masked inputs He et al. (2022); Tong et al. (2022); Wang et al. (2023b); Feichtenhofer et al. (2022), aligning videos with languages Li and Wang (2020); Xu et al. (2021); Yan et al. (2022); Li et al. (2023e), and predicting the next token using videos Alayrac et al. (2022); Sun et al. (2023c); Li et al. (2023d). These approaches have turned out to be complementary and can be unified through a progressive training scheme. Notably, methods such as InternVideo Wang et al. (2022), UMT Li et al. (2023e), and VideoPrism Zhao et al. (2024) have utilized a two-stage training approach involving masked reconstruction and multimodal contrastive learning, leading to improved performance in downstream tasks. Following this line, we aim to further extend this progressive training scheme by incorporating video-based next token prediction and scaling the entire training process, including models and data, to build a new family of video foundation models.
@@
The proposed video foundation model, coined as InternVideo2, is built through a progressive training scheme. The learning involves three stages: (1) capturing spatiotemporal structure via unmasked reconstruction, (2) aligning with semantics from other modalities, and (3) enhancing its open-ended dialogue power through next token prediction. In the initial stage, the model learns to reconstruct the unmasked video tokens, allowing the video encoder to develop basic spatiotemporal perception capability. To estimate the existing tokens, vision encoders (InternViT Chen et al. (2023c) and VideoMAE-g Wang et al. (2023b)) trained differently are employed as proxies. In the next stage of crossmodal learning, the architecture is expanded to include audio and text encoders. This not only improves the alignment between videos and text but also endows InternVideo2 the ability to handle video-audio tasks. By incorporating these additional modalities, the model’s understanding of videos is enriched and aligned with their semantics. Finally, in the next-token prediction stage, a video-centric dialogue system and the corresponding instruction-finetuning dataset are built to further tune the InternVideo2. By connecting InternVideo2 to LLMs, the video encoder is further updated through next-token prediction training, enhancing its ability for open-ended tasks such as VQA and video description.
@@
For the training of InternVideo2, we emphasize the spatiotemporal consistency and labeling quality in the data. We build a large-scale multimodal video-centric dataset consisting of 402M data entries, which includes 2M videos, 50M video-text pairs (from WebVid Bain et al. (2021) and InternVid Wang et al. (2023d)), 50M video-audio-speech-text pairs (InternVid2), and 300M image-text pairs. Specifically, for InternVid2, we segment videos into clips semantically and focus on recalibrating the clip descriptions using three modalities: audio, video, and speech. We first generate captions for these three modalities separately. Then individual captions are fused together to create a more comprehensive description, which will improve the model’s ability to comprehend and interpret the video accurately.
@@
We evaluate InternVideo2 across a wide range of video-related tasks. These tasks span from basic spatiotemporal perception, such as action recognition, to high-level reasoning tasks, such as long video or procedure-aware question-answering (QA), as given in Fig. 1. The results (in Sec. 5) demonstrate that InternVideo2 achieves the state-of-the-art performance on multiple tasks and is able to analyze and reason over sequences of actions. This top performance signifies its capability to effectively capture and understand video content. These empirical findings validate that InternVideo2 could serves as a general video encoder for future exploration in video understanding. In summary, our contributions to video understanding are as follows.
@@
This paper introduces InternVideo2, a competitive family of video foundation models that leverages masked reconstruction, crossmodal contrastive learning, and next token prediction to make model perceptive, semantic, and capable of reasoning in video understanding.
InternVideo2 achieves the state-of-the-art performance for more than 60 video / audio tasks. Our model demonstrates superior performance in video-related dialogue and long video understanding, highlighting its potential in modeling high-level world knowledge.
@@
We provide an enhanced dataset to train InternVideo2. This includes the validation and incorporation of audio data during training, as well as the improved captioning method. These improvements result in significant enhancements in model performance and generalization ability.
@@ 2 Related Work
Video Foundation Models. Studies on learning video foundation models becomes increasingly crucial considering its wide applications Li and Wang (2020); Xu et al. (2021); Li et al. (2023e); Wang et al. (2022); Zhao et al. (2024); Wang et al. (2023c); Yan et al. (2022); Feichtenhofer et al. (2022); Wang et al. (2023a); Tong et al. (2022); Wang et al. (2023b). Typical methods in building video foundation models (ViFM) include video-text contrastive learning Li and Wang (2020); Xu et al. (2021); Wang et al. (2022), masked video modeling Tong et al. (2022); Wang et al. (2023b, 2022); Fu et al. (2021), and next token prediction Alayrac et al. (2022); Sun et al. (2023c, a). Specifically, All-in-one Wang et al. (2023a) utilized a single backbone with unified multiple pretraining objectives. On the other hand, UMT Li et al. (2023e) combined masked modeling with video-text contrastive learning, demonstrating strong performance in both action recognition and video-language tasks. Another approach is mPLUG-2 Xu et al. (2023), which introduced a new design for modulating different modalities. It shared a common module across modalities to enhance relations while incorporating modality-specific modules for discrimination. In addition to video-text pretraining, researchers have also explored the use of audio information in videos to improve performance. MERLOT Reserve Zellers et al. (2022) learned video representations using a large-scale dataset of video-speech-transcript pairs. VALOR Chen et al. (2023b) employed independent encoders for video, audio, and text and trains a joint visual-audio-text representation. VAST Chen et al. (2024b) constructed an audio-visual-speech dataset and develops a multimodal backbone that excels in video-audio-related tasks. VideoPrism Zhao et al. (2024) combined video-text contrastive learning and video token reconstruction on a combination of public and proprietary videos, achieving leading results across various video tasks.
@@
Multimodal Large Language Models. With advances in large language models (LLMs) Devlin et al. (2018); Raffel et al. (2020); Brown et al. (2020), their multimodal versions (MLLMs) is becoming popular as it can handle open-world tasks. Seminal works like Flamingo Alayrac et al. (2022) showed outstanding zero/few-shots performances over a range of multimodal tasks Goyal et al. (2017b); Plummer et al. (2015); Xu et al. (2016); Marino et al. (2019). Public MLLMs Zhu et al. (2023b); Liu et al. (2023); Gong et al. (2023) such as LLaVA Liu et al. (2023) and InstructBLIP Dai et al. (2023) proposed to use visual instruction-tuning data to improve the visual dialogue ability. Some video-centric MLLMs have been proposed, such as VideoChat Li et al. (2023c), VideoChatGPT Maaz et al. (2023b) and Valley Luo et al. (2023), by using instruction data to connect video encoders to LLMs for open-world video understanding.
@@ 3 Method
We learn InternVideo2 in three stages, illustrated in Fig. 2. The stages include spatiotemporal token reconstruction, video-audio-speech-language contrastive learning, and connecting to a large language model (LLM) for joint training.
@@
The video encoder used in InternVideo2 follows the Vision Transformer (ViT) Dosovitskiy et al. (2020) and includes additional projection layers for distillation. Inspired by previous works Chen et al. (2023c); Yu et al. (2022), we introduce attention pooling to the ViT. For input videos, we sparsely sample 8 frames Wang et al. (2016) and perform a 1414 () spatial downsampling. These spatiotemporal tokens are then concatenated with a class token and combined with 3D position embeddings. The details of ViT-6B architecture are given in Supp.
3.1 Stage1: Reconstructing Unmasked Video Tokens
We exploit two expert models to guide the video encoder to conduct token-level reconstruction for unmasked areas. Specifically, we adopt InternVL-6B Chen et al. (2023c) and VideoMAEv2-g Wang et al. (2023b) to transfer unmasked knowledge via simple projection layers. When training, we input the full videos into different teachers and mask out 80% of the tokens frame by frame, under the semantic guidance of the multimodal model InternVL and motion-aware model VideoMAEv2. We only align the unmasked tokens, by minimizing their mean squared error (MSE) between student and teachers. The learning objective is to reconstruct the remaining tokens as:
where , , and are our video encoder, InternViT-6B Chen et al. (2023c), and ViT-g of VideoMAEv2, respectively. stands for the token index and is the corresponding token extracted by InternVideo2 for input video . is the normalization factor. and balance the influence between the employed models.
In our implementation, we randomly initialize the video encoder and then align its outputs from different layers (transformed by learnable multilayer perceptrons) to those of expert models. Specifically, we align: 1) the last layers of InternVL, 2) the last layers of VideoMAEv2, and 3) the final output token of InternVL. These alignments are made to the corresponding outputs of the video encoder using the norm. The different loss terms are simply summed for optimization. After pretraining, we drop those projection layers and only use the basic encoder. Compared with only using the multimodal model in UMT and VideoPrism, our strategy makes the vision encoder multimodal-friendly as well as enhances its temporal sensitivity for action modeling.
@@ 3.2 Stage 2: Aligning Video to Audio-Speech-Text
We exploit the correspondence between video and audio, speech, and text to encourage InternVideo2 to learn more semantics. In practice, InternVideo2 has a huge video encoder, and its employed audio and text encoders are relatively lightweight. The used audio encoder is a 12-layer transformer initialized with BEATs Chen et al. (2023a) (90M). It takes in audio features, which are 64-dimensional log Mel filterbank spectrograms using a 25ms Hamming window, transformed from 10-second-long clips, padding with zeros. For the text and speech encoders, we initialize the text encoder and multimodal decoder using BERT-Large Devlin et al. (2018). Specifically, we utilize the initial 19 layers of BERT-Large as the text encoder, with the subsequent 5 layers equipped with cross-attention layers serving as the multimodal decoder.
For pretraining objectives, we establish alignment across different modalities via text, including video, audio, image, and speech. We employ crossmodal contrastive and matching losses with masked language reconstruction loss as:
The employed and are standard loss from Cheng et al. (2022). Specifically, the crossmodal contrastive learning is given as:
@@
where computes the matching likelihood between and . denotes whether the given video and text are paired () or not ().
@@
To improve the training efficiency, we employ the masked learning strategy, aligning unmasked video tokens to tokens from other modalities first, then using full video tokens reconstruction shortly. Specifically, it consists of two steps as follows:
@@
We freeze the audio encoder and focus on aligning visual, audio, and text features. For pre-training, we use a comprehensive set of image, video, and audio-video data. The combinations of modalities used are represented as where each pair denotes the concatenated features from the respective modalities.
@@
We freeze the vision encoder to jointly align audio, visual, and text features. Post-pretraining is conducted using a smaller subset of image and video data (25M samples), along with the full set of audio (0.5M samples) and audio-video data (50M samples). Since the parameters of the largest ViT-6B model are frozen, we do not use masking strategies in this phase to ensure consistency with the inference process and to minimize any performance degradation in downstream tasks. The modality combinations used here are .